# Learning Source Biases: Multi-sourced Misspecifications and Consequences

### v2

@@ A.2 Proofs of the concentration of limiting posterior belief on KL minimizer
$\leq-\frac{1}{2}N\log\lambda_{\min}(\Sigma^{\prime})\leq N|\log\underline{\lambda}|/2$
when $k_{2}(<\lambda_{\max}(\Sigma^{\prime}))$ is large, where the second inequality uses the bound that $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}$ and the assumption that $\underline{\lambda}<1$. Together, we obtain the following:
$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in\Theta}\log l(x\mid b^{\prime},\Sigma^{\prime})\right]\leq\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\mathbf{1}_{\|\omega+\varepsilon\|\leq k_{1}/2}\sup_{(b^{\prime},\Sigma^{\prime})\in D}l(x\mid b^{\prime},\Sigma^{\prime})\right]+N|\log\underline{\lambda}|/2$
$\leq-\frac{1}{2}\left[N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right]\mathbb{P}(\|\omega+\varepsilon\|\leq k_{1}/2)+N|\log\underline{\lambda}|/2,$
@@
It remains to verify that $\Theta-D$ is compact. By definition, this set consists of all vectors $b^{\prime}$ such that $\|b^{\prime}-b^{*}\|\leq k_{1}$, together with all covariance matrices of the form $\Omega^{\prime}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, whereby the eigenvalues of $\Omega^{\prime}$ belong to $[\underline{\lambda},k_{2}]$. The last set is compact, and this completes the proof. ∎
@@
In the environment considered in the baseline model, the support of the DM’s prior belief is
$\Theta=\{(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\}$
where $\Omega^{*}$ denotes the true covariance matrix of Gaussian errors. Since $\Omega^{*}\in\Lambda$, all we need is to replace every $\Sigma$ and $\Sigma^{\prime}$ in the proof of Lemma 4 with $\Sigma^{*}\coloneqq\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. ∎
A.3 Proof of Lemma 2
Since $\Sigma^{*}$ is a positive definite matrix, it is invertible. Applying the block matrix inversion formula to $\Sigma^{*}$ yields
where $A$, $B$, $C$, and $D$ were defined in Section 3.3. The system (3) of first-order conditions associated with KL minimization, in its matrix form, is
Multiplying both sides of this equation by $D-CA^{-1}B$ yields $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ as desired.
Using this result to simplify $\Sigma^{*-1}\Delta$, we obtain the following:
$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}+A^{-1}B(D-CA^{-1}B)^{-1}(CA^{-1}\widetilde{\Delta}-\widehat{\Delta})\\ \hline\cr\mathbf{0}_{N-M}\end{array}\right]$ (14)
Simplifying $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-\hat{b})-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=-\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta$ accordingly yields $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$ as desired. ∎
A.4 Proofs of results that characterize KL minimizers
@@
The environment described in the proposition features $\Sigma^{*}=\text{diag}(v)+\alpha\alpha^{\top}$, $A=\text{diag}(v_{1},\cdots,v_{M})+[\alpha_{1},\cdots,\alpha_{M}]^{\top}[\alpha_{1},\cdots,\alpha_{M}]$ and $C=[\alpha_{M+1},\cdots,\alpha_{N}]^{\top}[\alpha_{1},\cdots,\alpha_{M}]$. Since $A$ is a rank-one update of an invertible matrix,
$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\begin{bmatrix}\alpha_{1}\nu_{1}\\ \vdots\\ \alpha_{M}\nu_{M}\end{bmatrix}\begin{bmatrix}\alpha_{1}\nu_{1}&\cdots&\alpha_{M}\nu_{M}\end{bmatrix}$
by the Sherman-Morrison formula. Substituting the expressions for $A^{-1}$ and $C$ into Lemma 2 yields
$\widehat{\Delta}=CA^{-1}\widetilde{\Delta}=\begin{bmatrix}\alpha_{M+1}\\ \vdots\\ \alpha_{N}\end{bmatrix}\left(1-\frac{\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\right)\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}=\begin{bmatrix}\alpha_{M+1}\\ \vdots\\ \alpha_{N}\end{bmatrix}\overline{\Delta},$
where $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i})$. To calculate the distortion in the best-linear predictor, multiply (17) by $\alpha^{\top}$ to obtain
$\alpha^{\top}\Sigma^{*-1}\Delta=[\alpha_{1},\cdots,\alpha_{M}]A^{-1}\widetilde{\Delta}=\left(1-\frac{\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\right)\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}=\overline{\Delta}.$
Since $\text{BLP}^{*}(X)=\alpha^{\top}\Sigma^{*-1}(X-b^{*})$ and $\widehat{\text{BLP}}(X)=\alpha^{\top}\Sigma^{*-1}(X-\hat{b})$ in the current environment, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\alpha^{\top}\Sigma^{*-1}\Delta=-\overline{\Delta}$ as desired. ∎
Proof of Proposition 2.
With a common latent factor, the KL minimization problem becomes
$\min_{\Delta\in\mathbb{R}^{N+1}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i\in\mathcal{M}}\frac{1}{2}\Delta^{\top}\Phi\Delta\text{ where }\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma.$
Since $\Sigma^{*}=\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, $\Sigma^{*-1}=\text{diag}(\nu)-\nu\nu^{\top}/(1+\mathbf{1}_{N}^{\top}\nu)$ by the Sherman-Morrison formula. Thus
$\Phi=\Gamma^{\top}\text{diag}(\nu)\Gamma-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\Gamma^{\top}\nu\nu^{\top}\Gamma,$
where the first part of the right-hand side equals (recall that $\Gamma=[\mathbb{I}_{N\times N}\mid\mathbf{1}_{N}]$)
$\left[\begin{array}[]{c}\mathbb{I}_{N\times N}\\ \hline\cr\mathbf{1}_{N}^{\top}\end{array}\right]\text{diag}(\nu)\left[\begin{array}[]{c|c}\mathbb{I}_{N\times N}&\mathbf{1}_{N}\end{array}\right]=\left[\begin{array}[]{c|c}\text{diag}(\nu)&\nu\\ \hline\cr\nu^{\top}&\mathbf{1}_{N}^{\top}\nu\end{array}\right],$
and the second part equals
$\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\left[\begin{array}[]{c}\mathbb{I}_{N\times N}\\ \hline\cr\mathbf{1}_{N}^{\top}\end{array}\right]\nu\nu^{\top}\left[\begin{array}[]{c|c}\mathbb{I}_{N\times N}&\mathbf{1}_{N}\end{array}\right].$
Subtracting the second part from the first yields
$\Phi=\left[\begin{array}[]{c|c}\text{diag}(\nu)-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu\nu^{\top}&\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu\\ \hline\cr\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu^{\top}&1-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\end{array}\right].$
The system of first-order conditions associated with KL minimization is
$E\begin{bmatrix}\Delta_{M+1}\\ \vdots\\ \Delta_{N+1}\end{bmatrix}=-F\begin{bmatrix}\Delta_{1}\\ \vdots\\ \Delta_{M}\end{bmatrix},\text{ where }$
$E\coloneqq\begin{bmatrix}\Phi_{M+1,M+1}&\cdots&\Phi_{M+1,N+1}\\ \vdots&\ddots&\vdots\\ \Phi_{N+1,M+1}&\cdots&\Phi_{N+1,N+1}\end{bmatrix}\text{ and }F\coloneqq\begin{bmatrix}\Phi_{M+1,1}&\cdots&\Phi_{M+1,M}\\ \vdots&\ddots&\vdots\\ \Phi_{N+1,1}&\cdots&\Phi_{N+1,M}\end{bmatrix}.$
From the expression for $\Phi$, we can read off those for $E$ and $F$:
$E=\text{diag}(\nu_{M+1}\cdots,\nu_{N},1)-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{M+1},&\cdots,&\nu_{N},&-1\end{bmatrix}$
and
$F=-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
Since $E$ is a rank-one update of an invertible matrix, applying the Sherman-Morrison formula shows that
$E^{-1}=\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)$
$+\frac{\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{M+1},\cdots,\nu_{N},-1\end{bmatrix}\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)/(1+\mathbf{1}_{N}^{\top}\nu)}{1-\begin{bmatrix}\nu_{M+1},\cdots,\nu_{N},-1\end{bmatrix}\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}/(1+\mathbf{1}_{N}^{\top}\nu)}$
$=\text{diag}(v_{M+1},\cdots,v_{N},1)+\frac{1}{\sum_{i=1}^{M}\nu_{i}}\left[\begin{array}[]{c}\mathbf{1}_{N-M}\\ \hline\cr-1\end{array}\right]\left[\begin{array}[]{c|c}\mathbf{1}_{N-M}^{\top}&-1\end{array}\right].$
Simplifying the system of first-order conditions accordingly yields
$\left[\begin{array}[]{c}\Delta_{M+1}\\ \vdots\\ \Delta_{N}\\ \hline\cr\Delta_{N+1}\end{array}\right]=\overline{\Delta}\left[\begin{array}[]{c}\mathbf{1}_{N-M}\\ \hline\cr-1\end{array}\right],\text{ where }\overline{\Delta}\coloneqq\frac{\sum_{i=1}^{M}\nu_{i}\Delta_{i}}{\sum_{i=1}^{M}\nu_{i}}.\qed$
@@
$\Sigma^{*}=\text{diag}\left(\Omega_{1},\cdots,\Omega_{N}\right)+\left(\mathbf{1}_{N}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{N}^{\top}\otimes\mathbb{I}_{K\times K}\right),$
where $\otimes$ denotes the Kronecker product. Write $\Sigma^{*}$ as
$\left[\begin{array}[]{c|c}A_{KM\times KM}&B_{KM\times K(N-M}\\ \hline\cr C_{K(N-M)\times KM}&D_{K(N-M)\times K(N-M)}\end{array}\right],$
@@
$A^{-1}$ $=G^{-1}+G^{-1}x\left(\Omega_{0}^{-1}+yG^{-1}x\right)^{-1}yG^{-1}$
$=\text{diag}(\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1})-\begin{bmatrix}\Omega_{1}^{-1}\\ \vdots\\ \Omega_{M}^{-1}\end{bmatrix}\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\begin{bmatrix}\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1}\end{bmatrix}$
where the first equality uses the Woodbury formula (Woodbury, 1950). Substituting the expressions for $A^{-1}$ and $C$ into Lemma 2 yields
@@
$\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=0}^{M}\Omega^{-1}_{i}=\Omega_{0}\Omega_{0}^{-1}=\mathbb{I}_{K\times K}$
@@
The distortion in the best linear predictor can be obtained likewise, as follows:

### v3

@@ A.2 Useful lemmas and their proofs
$\leq-\frac{1}{2}N\log\lambda_{\min}(\Sigma^{\prime})\leq N|\log\underline{\lambda}|/2,$
where the first inequality is obvious, and the second one exploits the fact that $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}$ and the assumption that $\underline{\lambda}<1$. Taken together, we obtain that
$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in\Theta}\log l(x\mid b^{\prime},\Sigma^{\prime})\right]\leq\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\mathbf{1}_{\|\omega+\epsilon\|\leq k_{1}/2}\sup_{(b^{\prime},\Sigma^{\prime})\in D}l(x\mid b^{\prime},\Sigma^{\prime})\right]+N|\log\underline{\lambda}|/2$
$\leq-\frac{1}{2}\left[N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right]\mathbb{P}(\|\omega+\epsilon\|\leq k_{1}/2)+N|\log\underline{\lambda}|/2,$
@@
It remains to verify that $\Theta-D$ is compact. By definition, this set consists of vectors $b^{\prime}$ such that $\|b^{\prime}-b^{*}\|\leq k_{1}$, paired with covariance matrices of the form $\Omega^{\prime}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$ such that the eigenvalues of $\Omega^{\prime}$ belong to $[\underline{\lambda},k_{2}]$. The last set is compact, and this completes the proof. ∎
@@
In the environment considered in Section 2, the support of the DM’s prior beliefs is
$\Theta=\{(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\},$
where $\Omega^{*}$ denotes the true covariance matrix of the errors. Since $\Omega^{*}\in\Lambda$, all that is needed is to replace every $\Sigma$ and $\Sigma^{\prime}$ in the proof of Lemma 2 with $\Sigma^{*}\coloneqq\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. ∎
A.3 Proofs of theorems and propositions
@@
We already defined and derived the constrained least squares estimator in Footnote 7. It remains to derive the best linear predictor of the state without the normality assumption.
Under the specification $X=b+\omega\mathbf{1}_{N}+\epsilon$ where $\omega$ and $X-b$ both have mean zero, the best linear predictor is given by a vector $\beta$ that minimizes the prediction squared-error $\mathbb{E}[\left(\omega-\beta^{\top}(X-b)\right)^{2}]$. Since
$\mathbb{E}[\left(\omega-\beta^{\top}(X-b)\right)^{2}]$
$=\mathbb{E}[\omega^{2}]+\mathbb{E}[\beta^{\top}(X-b)(X-b)^{\top}\beta]-2\mathbb{E}[\beta^{\top}(X-b)\omega]$
$=1+\beta^{\top}\Sigma^{*}\beta-2\beta^{\top}\mathbf{1}_{N},$
where the last line is minimized at $\beta=\Sigma^{*-1}\mathbf{1}_{N}$, the BLP is $\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b).$
Proof of Theorem 2.
We only prove Part (i) of the theorem. Part (ii) of the theorem has already been established in the main text. Part (iii) of the theorem follows from straightforward algebra.
To begin, recall the KL miminization problem considered in Case (III) of HKS:
where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$. The solution $(\hat{f},\widehat{\Sigma})$ to this problem is shown to satisfy two properties: First, $\Delta f_{i}\coloneqq\hat{f}_{i}-f_{i}^{*}$ as if the DM knows the true covariance matrix $\Sigma^{*}$. Second, $\widehat{\Sigma}=\Sigma^{*}+(M\Delta f)(M\Delta f)^{\top}$. The assumption that only a single entry of $f$ is misspecified plays no role in this result, verified by checking the proof of HKS line by line.
Letting $f=b$ and $M=\mathbb{I}_{N\times N}$ in the above argument leads us to conclude that, in the environment considered in Section 5.1, the solution to the KL minimization problem would be $\Delta=\hat{b}_{i}-b_{i}^{*}$ and $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$ had our DM been able to choose any SPD matrix whose minimum eigenvalue is bounded below by $\underline{\underline{\lambda}}$. The complication arises from the restriction that the covariance matrix must take the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, where $\Omega$ is SPD with its minimum eigenvalue bounded below by $\underline{\lambda}<1\wedge\lambda_{\min}(\Omega^{*})$. We are left to verify that the solution to the unconstrained problem satisfies this additional constraint. If so, then it must also solve the constrained optimization problem.
To verify, note that the solution to the unconstrained problem: $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, is attained by $\widehat{\Omega}=\Omega^{*}+\Delta\Delta^{\top}$. $\widehat{\Omega}$ is clearly SPD, and $\lambda_{\min}(\widehat{\Omega})\geq\lambda_{\min}(\Omega^{*})>\underline{\lambda}$ by Lemma 3. This completes the proof. ∎
@@
$\text{diag}\left(\Omega_{1},\cdots,\Omega_{N}\right)+\left(\mathbf{1}_{N}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{N}^{\top}\otimes\mathbb{I}_{K\times K}\right),$
where $\otimes$ denotes the Kronecker product. Write this matrix as
$\left[\begin{array}[]{c|c}A_{KM\times KM}&B_{KM\times K(N-M)}\\ \hline\cr C_{K(N-M)\times KM}&D_{K(N-M)\times K(N-M)}\end{array}\right],$
@@
$A^{-1}$ $=G^{-1}-G^{-1}x\left(\Omega_{0}^{-1}+yG^{-1}x\right)^{-1}yG^{-1}$
$=\text{diag}(\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1})-\begin{bmatrix}\Omega_{1}^{-1}\\ \vdots\\ \Omega_{M}^{-1}\end{bmatrix}\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\begin{bmatrix}\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1}\end{bmatrix},$
where the first equality uses the Woodbury formula (Woodbury, 1950). Substituting the expressions for $A^{-1}$ and $C$ into Theorem 1 yields
@@
$\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=0}^{M}\Omega^{-1}_{i}=\Omega_{0}\Omega_{0}^{-1}=\mathbb{I}_{K\times K},$
@@
Additionally,